Find the indicated dot product.

r = <9, -7, -8>, v = <3, 4, 7>, w = <6, -9, 7>

v ⋅ w

help me please
1
Expert's answer
2014-05-14T13:28:22-0400
Dot product v &sdot; w=3*6+4*(-9)+7*7=18-36+49=18+13=31 (it is calculated as the sum of the products of the corresponding entries of the two sequences of numbers).